OpenStudy (anonymous):

Would this series converge or diverge? sum from n=1 to infinity (1/(3n-2)^(n+1/2) )? What test would you use?

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} \frac{ 1 }{ (3n-2)^{n+\frac{ 1 }{ 2 }} }\]

OpenStudy (amistre64):

have you tried a ratio test yet?

OpenStudy (anonymous):

I didn't know what to do when I got \[\lim_{n \rightarrow \infty} \ \frac{ (3n-2)^{n+\frac{ 1 }{ 2 }} }{ (3n+1)^{n+\frac{ 3 }{ 2 }} }\] I'm thinking test for divergence where I treat this as a function, so I can just say it diverges?

OpenStudy (anonymous):

I think a comparison test might get you somewhere. I'll try it out.

OpenStudy (anonymous):

Looks like it does. Try comparing to \[\sum_{n=1}^\infty\frac{1}{(3n-2)^n}\] And use the root test to show convergence/divergence of the comparison series.

OpenStudy (anonymous):

Root test? lol our professor skipped that section, we only went over ratio.

OpenStudy (anonymous):

I'm pretty sure the ratio test also works.

OpenStudy (amistre64):

integral test seem to be tricky at best ....

OpenStudy (anonymous):

By the ratio test, you get a similar-looking limit: \[\lim_{n\to\infty}\frac{(3n-2)^n}{(3n+1)^{n+1}}\\ \lim_{n\to\infty}\frac{1}{3n+1}\left(\frac{3n-2}{3n+1}\right)^n\\ \]

OpenStudy (amistre64):

\[\sum_{n=1}^{\infty} \frac{ 1 }{ (3n-2)^{n+\frac{ 1 }{ 2 }} }\] \[\sum_{n=1}^{\infty} \frac{ 1 }{ \sqrt{3n-2} }\frac{ 1 }{ (3n-2)^{n} }\]

OpenStudy (anonymous):

^^^If you can figure out that limit, you can skip the comparison stuff.

OpenStudy (amistre64):

the wolf says it converges .... :)

OpenStudy (anonymous):

am I to use the ratio test on the function you separated? \[\lim_{n \rightarrow \infty} \frac{ (3n-2)^{1/2} }{ (3n+1)^{1/2} }\frac{ (3n-2)^{n} }{ (3n+3)^{n+1} }\] looks messy?

OpenStudy (amistre64):

\[\frac{ \sqrt{3n-2} }{ \sqrt{3n+1} }\frac{ (3n-2)^{n} }{ (3n+1)^{n+1} }\] \[\cancel{\sqrt{\frac{ {3n-2} }{ {3n+1}}}}^1\frac{ (3n-2)^{n} }{ (3n+1)^{n+1} }\] \[\cancel{\frac{ (3n)^n+....}{ (3n)^{n+1}+.... }}^0\text{ since the bottom is a higher degree than the top}\]

OpenStudy (amistre64):

so id say it does converge

OpenStudy (anonymous):

That's brilliant :) Thanks

OpenStudy (amistre64):

.. good luck ;)